The longer base of the isosceles trapezoid is 12 cm. The side is 5 cm. The acute angle is 37 degrees. Find the shorter base.
Please tell me the full answer, if it is 8.006 degrees

Answer by ikleyn(52781)

.
the shorter base = 12 - 2*5*cos(37°) = 12 - 2*5*0.799 = 4.01.